OpenStudy (anonymous):

How to solve this inequality: -23>=g-6

OpenStudy (anonymous):

Thank you :)

OpenStudy (akshay_budhkar):

My pleasure :D

OpenStudy (akshay_budhkar):

ok i made a blunder in my sleep :P

OpenStudy (akshay_budhkar):

it is

OpenStudy (akshay_budhkar):

\[-23 + 6 \ge g\] therefore \[-17 \ge g\] \[g \le -17\]

OpenStudy (akshay_budhkar):

sorry